OpenStudy (anonymous):

Use completing the square to solve (x + 7)(x – 9) = 25 for x.

OpenStudy (solomonzelman):

first expand the product. (x+7)(x-9) = ?

OpenStudy (anonymous):

x^2-9x+7x-63

OpenStudy (solomonzelman):

yes, (x + 7)(x – 9) = 25 x\(^2\)-9x+7x-63 = 25 add the x's on the left side, x\(^2\)-2x-63 = 25

OpenStudy (solomonzelman):

Add 63 to both sides, and u then get?

OpenStudy (anonymous):

-38

OpenStudy (solomonzelman):

x\(^2\)-2x-63 = 25 x\(^2\)-2x-63\(\small\color{red}{+63}\)= 25\(\small\color{red}{+63}\)

OpenStudy (solomonzelman):

like that, add 63 to both sides. what do you get after doing this?

OpenStudy (anonymous):

ooooooooooh i see. x^2-2x= 88

OpenStudy (solomonzelman):

yes, and then tell me what number would make the x²-2x +what a perfect square trinomial? (remember the rule DanJS posted: "half of the coefficient on x, then squared")

OpenStudy (anonymous):

1?

OpenStudy (solomonzelman):

that number that makes the left hand side a perfect square trinomial, this number, you should add to both sides of the equation.

OpenStudy (solomonzelman):

yes, it is 1.

OpenStudy (solomonzelman):

So, add 1 to both sides

OpenStudy (solomonzelman):

x\(^2\)-2x\(\small\color{blue}{+1}\)= 88\(\small\color{blue}{+1}\)

OpenStudy (solomonzelman):

can you go on from here? steps just in case: \(1.\) simplify the right side \(2.\) factor the left side (it should look like " (x-number)² ") \(3.\) take the square root of both sides (and don't forget the \(\pm\)) \(4.\) add 1 to both sides of the equation